# LEGO: Learning and Graph-Optimized Modular Tracker for Online Multi-Object Tracking with Point Clouds

### v3

@@ III-F Association Module
In 3D MOT, the data association module plays a crucial role in determining the correspondence between detected and predicted objects, to update the state of the objects. Most of the methods used in 3D MOT rely on the minimization of the geometry association cost matrix. This is done by evaluating the overlap or distance between the 3D bounding boxes of the predicted and detected objects, using the global nearest neighbor principle and typically the Hungarian algorithm [61]. Recently, [38] proposed to incorporate geometry, motion and appearance features from images to construct such association cost matrix. In our proposed approach, we follow the similar idea to define the geometry and motion cost matrix, but using GDAN score matrix $\bm{B}$ to replace the appearance features. The details are discussed in the follow parts.
@@
where, $w_{ang}$ is the weight for angle difference, $w_{velo}$ is the weight for velocity difference, whereas $\tilde{v}$ and $\hat{v}$ denote the velocities of the predicted object states and detected object states respectively.
However, the current approach for calculating matrix $\bm{A}$ is limited to considering only the geometry and motion costs, neglecting the valuable information embedded in the 3D features present within the 3D point cloud. To address this limitation, we propose the final association cost matrix $\bm{C}$, which incorporates the geometry and motion cost matrix $\bm{A}$ and the GDAN score matrix $\bm{B}$, respectively. This amalgamation is achieved through a weighted linear combination of $\bm{A}$ and $\bm{B}$:
@@
where $w_{B}$ represents the weight assigned to the GDAN score matrix $\bm{B}$. Notably, the subtraction operation between matrix $\bm{A}$ and $\bm{B}$ in Eq. (14) is performed due to the score being determined as a negative cost. By incorporating the GDAN score matrix alongside the geometry and motion cost matrix, our proposed method enriches the data association process, allowing for a more comprehensive evaluation of 3D features. With the final association cost matrix $\bm{C}$, as input, the Hungarian algorithm is used to solve the 2D assignment problem, identifying the matched pairs and unmatched objects.
@@ III-G State Update Module
Upon acquiring the association map from the association module, the Kalman filter is utilized to update the state of the predicted state. The updated mean $\bm{X}_{t}$ and covariance $\bm{P}_{t}$ are computed through the following Kalman update equations:
@@ III-H Track Management Module
In addition to the matched predicted states and the detected states, we should also deal with unmatched detected states, which may arise in two different cases: 1) Objects naturally disappear from the field of vision, and 2) Objects are misdetected, possibly due to temporary occlusion by other objects or movement beyond the optimal range of the LiDAR sensor.
@@ IV Experiments and Analysis
In this section, we will discuss the results of our experiment. We will begin by providing an overview of the dataset used, followed by a description of the implementation details. Next, we will present a thorough ablation study. Lastly, we will compare our results with that of the baseline models.
@@ IV-B Implementation Details
GDAN score calculation module: This module is characterized by setting the number of heads in each attention structure to 3, and GLUE is employed as the activation function. The association self-attention block incorporates softmax as the normalization function for the adjacency matrix.
@@ IV-C1 Quality of Detection Input
The performance of a tracker is inherently tied to the effectiveness of the integrated detector. In the conducted experiments, three distinct object detectors were assessed: the CasA detector, the PointGNN detector, and the PointRCNN detector. These detectors were chosen based on their performance in the KITTI 3D Object detection challenge and arranged in descending order of effectiveness. As depicted in Table I, the tracking performances align with the quality of the detection input for all the assessed trackers. The result of our LEGO has been visualized in Fig. 6. The first row and the second show the ground truth and tracking result reported by our LEGO in scene 0006, the detected state in the current frame is marked by a bounding box and its track ID. The third row and the fourth-row show results in scene 0008.
IV-C2 Performance Comparison with State-of-the-Art Trackers Using LiDAR Only
As delineated in Table I, a series of comparative analyses were conducted on various tracking methodologies within the context of the KITTI tracking benchmark. First, we compare our method with PC3T [38], utilizing PointRCNN as the object detector. The results reveal that our method achieved an enhancement of 0.25 in HOTA and 0.37 in MOTA on the testing dataset. Further analysis is extended to other methodologies utilizing disparate detectors, as outlined in Table I. For instance, in a comparison with CenterTube [34], which employs the same detector PointGNN, our method demonstrate an improvement of 8.32 in HOTA and 1.17 in MOTA. Likewise, when compared with UG3DMOT [36] that utilizes the CasA detector, our method achieved an enhancement of 2.15 in HOTA and 2.63 in MOTA.
IV-C3 Performance Comparison with State-of-the-Art Trackers Using LiDAR and Camera
@@ IV-C4 Comparison of Proposed Method and Other State-of-the-Art Trackers on 3D MOT Metrics
Our proposed LEGO tracker has been subjected to extensive comparative evaluation against various trackers, employing front-view 3D MOT metrics, such as sAMOTA, AMOTA, and AMOTP. The details of these comparisons are tabulated in Table III. Against GNN3DMOT [43], LEGO demonstrated marked improvement across several fronts, with enhancements of 1.22 in sAMOTA, 2.51 in AMOTA, and 8.87 in AMOTP. The comparisons were further extended to trackers utilizing PointGNN as their detector. When juxtaposed with CAMO-MOT [43], LEGO matched the SAMOTA, but exhibited improvement in AMOTA and AMOTP, by 0.06 and 5.57, respectively. Additionally, LEGO was evaluated against CenterTube, which utilizes CenterPoint as its detector, a system generally considered to offer superior detection performance over PointGNN. Nevertheless, as elucidated in Table III, despite using PointGNN as the detector, LEGO managed to secure improvements of 1.31 in sAMOTA, 1.86 in AMOTA, and 6.82 in AMOTP.
@@ IV-D1 Effectiveness of Weight between Matrix $\bm{A}$ and Matrix $\bm{B}$
Throughout the experimental phase, refinements were introduced to the cost weight $w_{B}$ to ascertain the optimal equilibrium between the geometry and motion cost matrix $\bm{A}$, and the GDAN score matrix $\bm{B}$. A comprehensive assessment was conducted across four distinct levels for $w_{B}$, specifically 0, 1, 2, 3, and 4, as described in Table IV. It was found that when $w_{2}$ was set to 2, the best outcomes were achieved. These results underscore the pivotal role played by the cost weight in determining the relative contributions of matrix $\bm{A}$ and matrix $\bm{B}$ to the association process.
@@ V Conclusion
In this paper, we propose an online tracker, LEGO, that effectively integrates the geometry and motion cost matrix and GDAN score matrix to tackle the inherent challenges of MOT. The proposed tracker leverages 3D point cloud information, topological structures, and motion prediction features to enhance tracking accuracy and robustness. The incorporation of message passing and attention mechanisms within the GDAN enables efficient estimation of future motion states and the effective utilization of 3D point cloud data. Furthermore, our offset correction module demonstrates the capability to rectify certain errors in the detection results, contributing to improved tracking performance. However, it is important to acknowledge that the use of graph structures in our approach may lead to longer training times due to increased computational demands. As a potential solution for future research, techniques, such as voxel downsampling could be implemented to reduce the parameter count in the GNN, thereby enhancing computational efficiency and reducing time consumption.

### v4

@@ III-F Association Module
In 3D MOT, the data association module plays a crucial role in determining the correspondence between detected and predicted objects, to update the state of the objects. Most of the methods used in 3D MOT rely on the minimization of the geometry association cost matrix. This is done by evaluating the overlap or distance between the 3D bounding boxes of the predicted and detected objects, using the global nearest neighbor principle and typically the Hungarian algorithm [70]. Recently, [44] proposed to incorporate geometry, motion and appearance features from images to construct such association cost matrix. In our proposed approach, we follow a similar idea to define the geometry and motion cost matrix, but using AMGN score matrix $\bm{B}$ to replace the appearance features. The details are discussed in the follow parts.
@@
where $w_{ang}$ is the weight for angle difference, $w_{velo}$ is the weight for velocity difference, whereas $\tilde{v}$ and $\hat{v}$ denote the velocities of the predicted object states and detected object states respectively.
However, the current approach to calculating the matrix $\bm{A}$ is limited to considering only geometry and motion costs, neglecting the valuable information embedded in the 3D features present within the 3D point cloud. To address this limitation, we propose the final association cost matrix $\bm{C}$, which incorporates the geometry and motion cost matrix $\bm{A}$ and the AMGN score matrix $\bm{B}$, respectively. This amalgamation is achieved through a weighted linear combination of $\bm{A}$ and $\bm{B}$:
@@
where $w_{B}$ represents the weight assigned to the AMGN score matrix $\bm{B}$. Notably, the subtraction operation between matrix $\bm{A}$ and $\bm{B}$ in Eq. (17) is performed due to the score being determined as a negative cost. By incorporating the AMGN score matrix alongside the geometry and motion cost matrix, our proposed method enriches the data association process, allowing for a more comprehensive evaluation of 3D features. With the final association cost matrix $\bm{C}$, as input, the Hungarian algorithm is used to solve the 2D assignment problem, identifying the matched pairs and unmatched objects.
@@ III-G State Update Module
Upon acquiring the association map from the association module, the Kalman filter is utilized to update the state of the predicted state, the same as in most of the existing works. The updated mean $\bm{X}_{t}$ and covariance $\bm{P}_{t}$ are computed through the following Kalman update equations:
@@ III-H Track Management Module
In addition to the matched predicted states and the detected states, we should also deal with unmatched detected states, which may arise in two different cases: 1). Objects naturally disappear from the field of vision, and 2). Objects are misdetected, possibly due to temporary occlusion by other objects or movement beyond the optimal range of the LiDAR sensor.
@@ IV Experiments and Analysis
In this section, we discuss the results of our experiment. We will begin by providing an overview of the dataset used, followed by a description of the implementation details. Next, we will present a thorough ablation study. Lastly, we will compare our results with those of the baseline models.
@@
    The metrics are reported by using 3D clustering approach rather than any learning-based model as 3D object detector, according to [39].
@@ IV-B Implementation Details
AMGN score calculation module: The number $K$ of AMGN module is set to 3.
@@ IV-C1 Quality of Detection Input
The performance of a tracker is inherently tied to the effectiveness of the integrated detector. In the conducted experiments, three distinct object detectors were assessed: the CasA 3D detector, the PointGNN 3D detector, and the PointRCNN 3D detector. These 3D detectors were chosen based on their performance in the KITTI 3D Object detection challenge and arranged in descending order of effectiveness.
IV-C2 Performance Comparison with State-of-the-Art Trackers Using LiDAR Only on 2D MOT Metrics
The typical tracking result of our LEGO has been visualized in Fig. 5. The first row and the second show the ground truth and tracking result reported by our LEGO in scene 0006, the track state for every object in the current frame is marked by a bounding box and its track ID. The third row and the fourth-row show results in scene 0008. As delineated in Table I, a series of comparative analyses were conducted on various tracking methodologies within the context of the KITTI tracking benchmark. First, we compare our method with PC3T [44], utilizing PointRCNN as the object detector. The results reveal that our method achieved an enhancement of 0.25 in HOTA and 0.37 in MOTA on the testing dataset. Further analysis is extended to other methodologies utilizing disparate detectors, as outlined in Table I. For instance, in a comparison with CenterTube [40], which employs the same detector PointGNN, our method demonstrate an improvement of 8.32 in HOTA and 1.17 in MOTA. Likewise, when compared with UG3DMOT [42] that utilizes the CasA detector, our method achieved an enhancement of 2.15 in HOTA and 2.63 in MOTA.
IV-C3 Performance Comparison with State-of-the-Art Trackers Using LiDAR and Camera on 2D MOT Metrics
@@ IV-C4 Comparison of Proposed Method and Other State-of-the-Art Trackers on 3D MOT Metrics
Our proposed LEGO tracker has also been subjected to extensive comparative evaluation against various trackers for tracking in 3D space, by employing 3D MOT metrics, such as sAMOTA, AMOTA, and AMOTP. The details of these comparisons are tabulated in Table III. By employing PointRCNN as 3D detector, LEGO demonstrated the best overall performance across several fronts, comparing to all other recent proposed trackers using LiDAR only. Additionally, LEGO was evaluated against CenterTube [40] which utilizes CenterPoint as its 3D detector, generally considered to offer superior detection performance over PointGNN. Despite using PointGNN as the 3D detector, LEGO still managed to secure improvements of 1.31 in sAMOTA, 1.86 in AMOTA, and 6.82 in AMOTP. The comparisons were further extended to trackers with fusion of LiDAR and camera. Utilizing PointGNN as 3D detector, when juxtaposed with CAMO-MOT [5], LEGO matched the sAMOTA, but exhibited improvement in AMOTA and AMOTP, by 0.06 and 5.57, respectively. Besides, the LEGO system also demonstrates significant enhancements when compared with the recently proposed approach [61], achieving a 7.15 improvement in AMOTP and a 5.67 increase in MOTA.
@@ IV-D1 Effectiveness of Weight between Matrix $\bm{A}$ and Matrix $\bm{B}$
Throughout the experimental phase, refinements were introduced to the cost weight $w_{B}$ to ascertain the optimal equilibrium between the geometry and motion cost matrix $\bm{A}$, and the AMGN score matrix $\bm{B}$. A comprehensive assessment was conducted across four distinct levels for $w_{B}$, specifically 0, 1, 2, 3, and 4, as described in Table IV. It was found that when $w_{2}$ was set to 2, the best outcomes were achieved. These results underscore the pivotal role played by the cost weight in determining the relative contributions of matrix $\bm{A}$ and matrix $\bm{B}$ to the association process.
@@ V Conclusion
In this paper, we propose an online LiDAR-based tracker, LEGO, that introduces the offset correction module and AMGN modules, and effectively integrates them into the existing tracking framework to tackle the inherent challenges of MOT. Our offset correction module demonstrates the capability to rectify certain errors in the detection results, contributing to improved tracking performance. Furthermore, the AMGN modules which use the learnable adjacency matrix to estimate the association relationships between detected objects and predicted objects in an efficient way. However, it is important to acknowledge that the use of graph structures in our approach may lead to longer training times due to increased computational demands. As a potential solution for future research, techniques, such as voxel downsampling could be implemented to reduce the parameter count in the GNN, thereby enhancing computational efficiency and reducing time consumption.